# An alternative condition for the solvability of the Dirichlet problem for the minimal surface equation on non-mean convex domains

### v1

@@ Abstract
We propose an alternative condition for the solvability of the Diri-chlet problem for the minimal surface equation that applies to non-mean convex domains. This condition is derived from a second-order ordinary differential equation whose solution produces a barrier that appears to be novel in the context of barrier constructions. It admits an explicit formulation and, in the setting of Hadamard manifolds, reveals a direct and transparent relationship between the geometry of the domain and the behavior of the boundary data required for solvability. The condition also extends naturally to unbounded domains.
In the Euclidean case, it is not only more practical to verify but also less restrictive than the classical Jenkins - Serrin criterion, ensuring the existence of solutions in situations where that approach fail. Furthermore, unlike the Jenkins-Serrin condition, our appproach separates the geometric properties of the domain from its boundary data, providing a clearer and more manageable framework for solvability analysis.
@@ 1 Introduction
In [8], $B$ is constructed as follows. Since $\partial\Omega$ is compact and of class $C^{2}$, there exists $l>0$ (uniform for all $p\in\partial\Omega$) such that the intersection $B_{l}(p)\cap\partial\Omega$ consists of a single connected component that is the graph over a domain $D$ in a hyperplane parallel to $T_{p}\partial\Omega$ of a function $g\in C^{2}(D)$ satisfying $|\nabla g|<1$. Set $l^{\prime}=l/\sqrt{2}$. Also, since $\partial\Omega$ is compact and $C^{2}$, its principal curvatures are uniformly bounded; let $k$ be an absolute bound for these curvatures. Define
@@
In [3], $B$ is defined through an expression involving $|A|$ and a distance $\xi^{\prime}$ to $\partial\Omega$, $0<\xi^{\prime}\leq d_{0}$, being $B=+\infty$ at the mean convex points, where $d_{0}$ is such that the normal exponential map
@@
Given the above, we propose an alternative criterion to determine whether the Dirichlet problem (2) admits a solution for a given $\varphi\in C^{2}(\overline{\Omega})$. Our goal is to retain the geometric insights of previous results, while providing a more practical condition, at least in the Euclidean case, that also applies to unbounded domains and, in the case of Hadamard manifolds, can be phrased in a way that makes explicit the relationship between the geometry of the domain and the behavior of the boundary data that guarantees solvability (see Corollary 4). We are inspired by [10], which uses assumptions relating the $C^{2}$-norm of $\varphi$ and $\omega_{\varphi}(\overline{\Omega})$ with the radius $r$ of the circles that, at each $p\in\partial\Omega$, is tangent to $\partial\Omega$ at $p$ and lies in $\mathbb{R}^{2}\backslash\Omega$ (exterior circle condition). Our criterion arises from solving a second-order ODE (29), which appears to be a novel method for constructing barriers for the Dirichlet problem (2).
@@
In order to state our results, we introduce some notation. Let $i:M\rightarrow[0,\infty]$ denote the injectivity radius function of $M$.
@@
Therefore, our condition is less restrictive than that of [10], where $r\geq 1$ is required and, moreover, holds for arbitrary dimension. It also allows us to conclude the solvability of the Dirichlet problem (2) in examples where the Jenkins-Serrin condition cannot be applied, as illustrated at the end of Section 3.
Regarding the connection with the Jenkins–Serrin condition, we emphasize that, even though our argument depends on an extension of $\varphi\in C^{2}(\partial\Omega)$ into $\overline{\Omega}$, it is always possible to construct such an extension - still denoted by $\varphi$ - on a neighborhood $\Omega_{\varepsilon}\subset\Omega$ of $\partial\Omega$, for some $\varepsilon>0$, in such a way that $\sup_{\overline{\Omega}_{\varepsilon}}\left|\nabla\varphi\right|=\sup_{\partial\Omega}\left|D\varphi\right|$ and $\sup_{\overline{\Omega}_{\varepsilon}}\left|\nabla^{2}\varphi\right|=\sup_{\partial\Omega}\left|D^{2}\varphi\right|$, where here $D\varphi$ and $D^{2}\varphi$ denote the first and second tangential derivatives of $\varphi$ on $\partial\Omega$. (See, for example, the extension constructed in Section 2.1 of [3].)
@@ 1.1 Basic context
Proof. Straightforward calculation.
@@
Proof. A rather extensive computation yields
$\nabla_{\nabla w}\nabla w$ $=\nabla_{\nabla\varphi}\nabla\varphi+\psi^{\prime}(d)\left[\nabla_{E_{n}}\nabla\varphi+\nabla_{\nabla\varphi}E_{n}+\psi^{\prime\prime}(d)E_{n}\right]+\psi^{\prime\prime}(d)E_{n}(\varphi)E_{n}.$
Now, observing that $\left\langle\nabla_{\nabla\varphi}E_{n},E_{n}\right\rangle=0$, we conclude that
$\left\langle\nabla_{\nabla w}\nabla w,\nabla w\right\rangle$ $=\nabla^{2}\varphi(\nabla\varphi,\nabla\varphi)+\psi^{\prime}(d)\left[2\nabla^{2}\varphi(E_{n},\nabla\varphi)+\nabla^{2}d(\nabla\varphi,\nabla\varphi)\right]$
$\quad+\left[\psi^{\prime}(d)\right]^{2}\nabla^{2}\varphi(E_{n},E_{n})+\psi^{\prime\prime}(d)\left(E_{n}(\varphi)+\psi^{\prime}(d)\right)^{2},$
which proves the result.
@@
$+\left[\psi^{\prime}\left(d\right)\right]^{2}\tau-\psi^{\prime\prime}\left(d\right)\left(E_{n}\left(\varphi\right)+\psi^{\prime}\left(d\right)\right)^{2}.$ (24)
Now compute
@@ 1.2 Proofs of Theorem 3 and Corollary 4
then we can construct an upper barrier at each $p\in\partial\Omega^{-}$ by setting
$v_{p}(x)=\begin{cases}w_{p}(x),&\text{if }x\in\Lambda_{p,\delta},\\ \displaystyle\inf_{\partial\Lambda_{p,\delta}\setminus\partial\Omega}w_{p},&\text{if }x\in\Omega\setminus\Lambda_{p,\delta}.\end{cases}$
For a lower barrier at $p\in\partial\Omega^{-}$, we define $\xi_{p}:=\varphi-\psi\circ d$.
@@
$u\left(x\right)=\sup\left\{s\left(x\right);s\in C^{0}\left(\Omega\right)\text{ is generalized solution of }M\text{ and }s|_{\partial\Omega}\leq\varphi|_{\partial\Omega}\right\}.$
Since every subsolution used to define $u$ is bounded above by $\sup_{\Omega}\varphi$, the function $u$ is well-defined. Moreover, by Perron’s method, we have $u\in C^{\infty}\left(\Omega\right)$ and $M\left(u\right)=0.$ The barriers at the boundary ensures that $u\in C^{2}\left(\overline{\Omega}\right)$ and $u|_{\partial\Omega}=\varphi.$

### v3

# An alternative solvability criterion for the Dirichlet problem for the minimal surface equation and an application to the mean curvature flow

@@ Abstract
We propose an alternative condition for the solvability of the Dirichlet problem for the minimal surface equation that applies to non-mean convex domains. This condition is derived from a second-order ordinary differential equation whose solution produces a barrier that appears to be novel in the context of barrier constructions. It admits an explicit formulation and, in the setting of Hadamard manifolds, reveals a direct and transparent relationship between the geometry of the domain and the behavior of the boundary data required for solvability. The condition also extends naturally to unbounded domains.
In the Euclidean case, it is not only more practical to verify but also less restrictive than the classical Jenkins–Serrin criterion, ensuring the existence of solutions in situations where that approach fails. Furthermore, unlike the Jenkins–Serrin condition, our approach separates the domain’s geometric properties from its boundary data, providing a clearer and more manageable framework for solvability analysis.
Beyond the stationary theory, the same ODE-generated barrier also yields natural sub- and supersolutions for the mean curvature flow of graphs, allowing one to initiate and control the short-time evolution from non-mean convex boundaries. As a consequence, our barrier construction extends a classical result of Ecker–Huisken on short-time existence of graphical mean curvature flow to domains with non-mean convex boundary; in the mean convex setting, their original result is recovered.
@@ 1 Introduction
In [11], $B$ is constructed as follows. Since $\partial\Omega$ is compact and of class $C^{2}$, there exists $l>0$ (uniform for all $p\in\partial\Omega$) such that the intersection $B_{l}(p)\cap\partial\Omega$ consists of a single connected component that is the graph over a domain $D$ in a hyperplane parallel to $T_{p}\partial\Omega$ of a function $g\in C^{2}(D)$ satisfying $|\nabla g|<1$. Set $l^{\prime}=l/\sqrt{2}$. Also, since $\partial\Omega$ is compact and $C^{2}$, its principal curvatures are uniformly bounded; let $k$ be an upper bound. Define
@@
In [3], $B$ is defined through an expression involving $|A|$ and a distance $\xi^{\prime}\in(0,d_{0}]$, where $d_{0}$ is such that the normal exponential map
@@
Given the above, we propose an alternative criterion to determine whether the Dirichlet problem (2) admits a solution for a given $\varphi\in C^{2}(\overline{\Omega})$. Our goal is to retain the geometric insights of previous results, while providing a more practical condition at least in the Euclidean case that also applies to unbounded domains and, in the case of Hadamard manifolds, can be phrased in a way that makes explicit the relationship between the geometry of the domain and the behavior of the boundary data that guarantees solvability (see Corollary 4). We are inspired by [14], which uses assumptions relating the $C^{2}$-norm of $\varphi$ and $\omega_{\varphi}(\overline{\Omega})$ with the maximal radius of exterior tangent circles to $\partial\Omega$. Our criterion arises from solving a second-order ODE (29), which appears to be a novel method for constructing barriers for the Dirichlet problem (2).
@@
The ideas developed here extend naturally to the parabolic setting of graphical mean curvature flow. Consider the evolution
$\partial_{t}u=\mathcal{M}(u),$
where $\mathcal{M}$ is given by (1). The same differential inequality underlying the ODE that generates our barrier reappears in the derivation of a priori height and gradient controls for the parabolic problem. As a consequence, the structural conditions ensuring solvability of the stationary Dirichlet problem provide, at the same time, sufficient hypotheses to preserve graphicality and boundary regularity for short time along the flow, even when $\partial\Omega$ is not mean convex (see Section 2). In particular, our criterion yields a short-time preservation mechanism for graphical mean curvature flow that extends Theorem 1.1 of Ecker-Huisken [6] to the non-mean-convex setting, while recovering their result in the mean-convex case. This yields a bridge between the elliptic theory and the mean curvature flow of graphs and furnishes explicit sub/supersolutions usable at the parabolic level (see, e.g., [6, 4]).
In order to state our next results, we introduce some notation. Let $i:\nolinebreak M\rightarrow[0,\infty]$ denote the injectivity radius function of $M$.
@@
Therefore, our condition is less restrictive than that of [14], where $r\geq 1$ is required and, moreover, holds for arbitrary dimension. It also allows us to conclude the solvability of the Dirichlet problem (2) in examples where the Jenkins-Serrin condition cannot be applied, as illustrated at the end of Section 1.2.
Regarding the connection with the Jenkins–Serrin condition, we emphasize that, even though our argument depends on an extension of $\varphi\in C^{2}(\partial\Omega)$ into $\overline{\Omega}$, it is always possible to construct such an extension—still denoted by $\varphi$—on a neighborhood $\Omega_{\varepsilon}\subset\Omega$ of $\partial\Omega$, for some $\varepsilon>0$, in such a way that
$\sup_{\overline{\Omega}_{\varepsilon}}|\nabla\varphi|=\sup_{\partial\Omega}|D\varphi|\quad\text{and}\quad\sup_{\overline{\Omega}_{\varepsilon}}|\nabla^{2}\varphi|=\sup_{\partial\Omega}|D^{2}\varphi|,$
where $D\varphi$ and $D^{2}\varphi$ denote the first and second tangential derivatives of $\varphi$ on $\partial\Omega$ (see, for example, the extension constructed in Section 2.1 of [3]).
@@ 1.1 Basic context
Proof. Let $E_{n}=\nabla d$ and $\{E_{1},\dots,E_{n-1}\}$ tangent to the level set $\Sigma_{p,s}$, extended off $\Sigma_{p,s}$ by parallel transport along the radial geodesics from $p_{0}$. Since $|\nabla d|=1$ on the regular set of $d$, we have
$0=\tfrac{1}{2}E_{i}\!\left(|\nabla d|^{2}\right)=\langle\nabla_{E_{i}}\nabla d,\nabla d\rangle=\nabla^{2}d(E_{i},E_{n})$
for every $i=1,\dots,n$, proving the first assertion.
For $i,j\leq n-1$, $\nabla^{2}d(E_{i},E_{j})=\langle\nabla_{E_{i}}E_{n},E_{j}\rangle$. Decompose $\nabla_{E_{i}}E_{n}$ into tangential and normal parts relative to $\Sigma_{p,s}$: $\nabla_{E_{i}}E_{n}=(\nabla_{E_{i}}E_{n})^{\top}+(\nabla_{E_{i}}E_{n})^{\perp}$. But $E_{n}$ is the unit normal to $\Sigma_{p,s}$, thus the Weingarten map $A_{E_{n}}$ is defined by $A_{E_{n}}(E_{i})=-(\nabla_{E_{i}}E_{n})^{\top}$ and the second fundamental form by $II_{E_{n}}(E_{i},E_{j})=\langle A_{E_{n}}(E_{i}),E_{j}\rangle$. Hence
$\nabla^{2}d(E_{i},E_{j})=\langle\nabla_{E_{i}}E_{n},E_{j}\rangle=\left\langle(\nabla_{E_{i}}E_{n})^{\top},E_{j}\right\rangle=-\,II_{E_{n}}(E_{i},E_{j}),$
which yields the stated identity with $H_{E_{n}}(E_{i},E_{j}):=\langle-(\nabla_{E_{i}}E_{n})^{\top},E_{j}\rangle$. Taking the trace over the tangential directions,
$\Delta d=\sum_{k=1}^{n}\nabla^{2}d(E_{k},E_{k})=\sum_{i=1}^{n-1}\nabla^{2}d(E_{i},E_{i})+\nabla^{2}d(E_{n},E_{n})=-\sum_{i=1}^{n-1}II_{E_{n}}(E_{i}),$
since $\nabla^{2}d(E_{n},E_{n})=\langle\nabla_{E_{n}}E_{n},E_{n}\rangle=0$ by $|\nabla d|=1$ and metric-compatibility.
@@
Proof. We compute $\nabla_{\nabla w}\nabla w$ using linearity and the chain rule. Since $w=\varphi+\psi(d)$ and $\nabla(\psi\circ d)=\psi^{\prime}(d)\nabla d=\psi^{\prime}(d)E_{n}$,
$\nabla_{\nabla w}\nabla w=\nabla_{\nabla\varphi}\nabla\varphi+\psi^{\prime}(d)\left[\nabla_{\nabla\varphi}E_{n}+\nabla_{E_{n}}\nabla\varphi+\psi^{\prime\prime}(d)E_{n}\right]+\psi^{\prime\prime}(d)E_{n}(\varphi)E_{n}.$
Taking the inner product with $\nabla w=\nabla\varphi+\psi^{\prime}(d)E_{n}$ and noting $\langle\nabla_{\nabla\varphi}E_{n},E_{n}\rangle=\nolinebreak 0$ (since $\langle E_{n},E_{n}\rangle=1$), we obtain
$\langle\nabla_{\nabla w}\nabla w,\nabla w\rangle$ $=\nabla^{2}\varphi(\nabla\varphi,\nabla\varphi)+\psi^{\prime}(d)\big[2\,\nabla^{2}\varphi(E_{n},\nabla\varphi)+\nabla^{2}d(\nabla\varphi,\nabla\varphi)\big]$
$\quad+\big[\psi^{\prime}(d)\big]^{2}\nabla^{2}\varphi(E_{n},E_{n})+\psi^{\prime\prime}(d)\big(E_{n}(\varphi)+\psi^{\prime}(d)\big)^{2},$
which is (22) after factoring $|\nabla\varphi|$ where appropriate.
@@
$+\left[\psi^{\prime}\left(d\right)\right]^{2}\tau-\psi^{\prime\prime}\left(d\right)\left(E_{n}\left(\varphi\right)+\psi^{\prime}\left(d\right)\right)^{2}$ (24)
is non positive. Now compute
@@ 1.2 Proofs of Theorem 3 and Corollary 4
(which is the expression (17) since that $\sigma=\varrho(1-\theta)$), then we can construct an upper barrier at each $p\in\partial\Omega^{-}$ by setting
$v_{p}(x)=\begin{cases}\min\left\{w_{p}(x),\inf_{\partial\Lambda_{p,\delta}\setminus\partial\Omega}w_{p}\right\}&\text{if }x\in\Lambda_{p,\delta},\\ \inf_{\partial\Lambda_{p,\delta}\setminus\partial\Omega}w_{p},&\text{if }x\in\Omega\setminus\Lambda_{p,\delta}.\end{cases}$
For a lower barrier at $p\in\partial\Omega^{-}$, just work with $\xi_{p}:=\varphi-\psi\circ d$.
@@
$u\left(x\right)=\sup\left\{s\left(x\right);s\in C^{0}\left(\Omega\right)\text{ is generalized subsolution of }\mathcal{M}\text{ and }s|_{\partial\Omega}\leq\varphi|_{\partial\Omega}\right\}.$
Since every subsolution used to define $u$ is bounded above by $\sup_{\Omega}\varphi$, the function $u$ is well-defined. Moreover, by Perron’s method, we have $u\in\nolinebreak C^{\infty}\left(\Omega\right)$ and $\mathcal{M}\left(u\right)=0.$ The barriers at the boundary ensures that $u\in C^{2}\left(\overline{\Omega}\right)$ and $u|_{\partial\Omega}=\varphi.$
@@
2 Mean Curvature Flow of Graphs without Mean Convexity
The differential inequality leading to the ordinary differential equation
$\psi^{\prime\prime}-\sigma\,\psi^{\prime}+\varrho=0,$ (38)
introduced in the previous section, naturally reappears in the study of the mean curvature flow of graphs. The same function $\psi$ that provides elliptic barriers for the Dirichlet problem also yields parabolic barriers that allow one to start and control the evolution even when $\partial\Omega$ is not mean convex. Thus equation (38) plays a central role in connecting the elliptic and parabolic regimes.
In the elliptic construction, the auxiliary function is written in terms of the distance $d_{p}(x)=\mathrm{dist}(x,\Sigma_{p})$ to the exterior geodesic sphere $\Sigma_{p}$ associated with each point $p\in\partial\Omega^{-}$, since these spheres encode the geometric rr–exterior sphere condition.
In the parabolic setting, in order to obtain barriers that are compatible with the Dirichlet condition along the whole boundary, it is more natural to work with the distance to the boundary,
$d(x)=\mathrm{dist}(x,\partial\Omega).$
Although $d(x)$ and $d_{p}(x)$ need not to coincide as functions in a neighborhood of $p$, along $\partial\Omega$ the two functions have the same first and second derivatives in the normal direction at $p$, that is, their first and second jets at $p$ coincide. Since all the differential inequalities leading to the ODE (38) depend only on the values of $\nabla d$ and $\nabla^{2}d$ along the boundary, this second–order agreement is sufficient to transfer the elliptic barrier construction to the parabolic framework without any modification of the resulting ODE.
Let $\Omega\subset M^{n}$ be a $C^{2}$ domain in a complete Riemannian manifold and let $\varphi\in C^{2}(\overline{\Omega})$ satisfy the structural condition of Theorem 3 (or, in the Hadamard case, condition (19)). We conclude from the above arguments that there is a tubular neighborhood of $\partial\Omega$,
$U_{\delta}:=\{x\in\Omega:d(x)<\delta\},$
where the signed distance function $d$ to $\partial\Omega$ is of class $C^{2}$ and, chosen $\psi$ as in Section 1.1 (Proposition 7),
$w^{\pm}(x):=\varphi(x)\pm\psi(d(x)),\qquad x\in U_{\delta},$
are such that $w^{-}$ is a subsolution and $w^{+}$ is a supersolution for the operator $\mathcal{M}$ in $U_{\delta}$.
Theorem 10
Let $\Omega\subset M^{n}$ be a $C^{2}$ domain in a complete Riemannian manifold, and let $\varphi\in C^{2}(\overline{\Omega})$ satisfy the structural condition of Theorem 3 (or, in the Hadamard case, condition (19)). Let $u_{0}\in C^{2}(\overline{\Omega})$ satisfy
$u_{0}|_{\partial\Omega}=\varphi,~w^{-}(x)\leq u_{0}(x)\leq w^{+}(x)\text{ for all }x\in U_{\delta},\text{ }\left|u_{0}\right|\leq\inf_{\partial U_{\delta}\backslash\partial\Omega}w^{+}\text{,}$
where $w^{\pm}$ and $U_{\delta}$ are as defined above. Then there exists $T>0$, depending only on $\varphi$ (through $\tau(\varphi)$) and on the geometric parameters $(\lambda_{r},\mu_{r})$ arising in the construction of the boundary barriers and, in particular, independent of the choice of $u_{0}$, such that the initial–boundary value problem
$\begin{cases}\partial_{t}u=\mathcal{M}(u),&(x,t)\in\Omega\times(0,T),\\ u(x,0)=u_{0}(x),&x\in\Omega,\\ u(x,t)=\varphi(x),&(x,t)\in\partial\Omega\times[0,T),\end{cases}$ (39)
admits a classical solution
$u\in C^{2,1}(\overline{\Omega}\times[0,T]).$
Moreover, for every $t\in[0,T]$, the hypersurface
$\Gamma_{t}=\{(x,u(x,t)):x\in\Omega\}$
remains a graph over $\Omega$, even if $\partial\Omega$ is not mean convex.
Proof. The proof relies directly on the equation (38). Let $d(x)$ denote the signed distance to $\partial\Omega$ and consider $w^{\pm}$ and $U_{\delta}$ are as defined above.
Since $w^{\pm}$ are independent of $t$, we have $\partial_{t}w^{\pm}=0$, and because $\psi$ satisfies (38), the functions $w^{+}$ and $w^{-}$ verify
$\partial_{t}w^{+}-\mathcal{M}(w^{+})\geq 0,\qquad\partial_{t}w^{-}-\mathcal{M}(w^{-})\leq 0$
in the tubular neighborhood $U_{\delta}$ of $\partial\Omega$, and both coincide with $\varphi$ on $\partial\Omega$.
Under the quantitative restriction (17) (or (19)), we have $w^{-}(x)\leq\varphi(x)\leq w^{+}(x)$ on $\partial\Omega$ and $w^{-}\leq w^{+}$ throughout $U_{\delta}$. To apply the parabolic maximum principle globally, we extend $w^{\pm}$ from $U_{\delta}$ to whole $\Omega$ by a standard $C^{2}$ gluing argument, by smoothing the capped function
$W^{+}\left(x\right)=\left\{\begin{array}[]{c}\min\{w^{+}\left(x\right),\inf_{\partial\Omega_{\delta}\backslash\partial\Omega}w^{+}\}\text{ se }x\in\overline{\Omega}_{\delta}\\ \inf_{\partial\Omega_{\delta}\backslash\partial\Omega}w^{+}\text{ se }x\in\Omega\backslash\Omega_{\delta}\end{array}\right$.
(which is a $C^{0}$ supersolution in $\overline{\Omega}$ for the operator $\mathcal{M}$) in such a way that the supersolution property is preserved, and we call it again $W^{+}$ (analogous construction for $W^{-}$). From our hypothesis on $u_{0}$, it follows that $W^{-}\leq u_{0}\leq W^{+}$ on $\overline{\Omega}$ with $u_{0}=\varphi$ on $\partial\Omega$. The classical theory for quasilinear parabolic equations (see [13]) guarantees local existence and uniqueness provided the structure is uniformly parabolic. We thus verify this uniform parabolicity within the region bounded by $W^{-}$ and $W^{+}$.
Since $\psi$ is smooth and monotone, the derivatives $|\nabla w^{\pm}|$ are bounded by constants depending only on $\tau(\varphi)$ and $(\lambda_{r},\mu_{r})$. By the parabolic maximum principle, the solution $u$ of (39) remains confined between $W^{-}$ and $W^{+}$, so that
$W^{-}(x)\leq u(x,t)\leq W^{+}(x)\quad\text{for all }(x,t)\in\overline{\Omega}\times[0,T].$
Hence $|\nabla u|$ is uniformly bounded because $\psi^{\prime}(d)$ is bounded, and therefore the coefficients
$a^{ij}(\nabla u)=\delta^{ij}(1+|\nabla u|^{2})-u_{i}u_{j}$
remain positive definite. The equation (39) is thus uniformly parabolic during the time interval considered.
With this structure established, the fixed–point argument applies: one linearizes (39) around $u_{0}$, solves the linearized problem via the Schauder theory, and iterates. The limit yields a solution $u\in C^{2,1}$ on $[0,T]$, where $T$ depends only on the upper bound of $|\nabla u|$ obtained from the barriers, hence only on $\tau(\varphi)$, $\lambda_{r}$ and $\mu_{r}$.
The boundary control provided by $\psi$ replaces the classical mean convexity assumption. Since the gradient is uniformly bounded, the graph cannot develop vertical tangents and remains graphical for all $t\in[0,T]$. Interior gradient and curvature estimates, identical to those derived by Ecker–Huisken [6], ensure that no curvature blow–up occurs before $T$. Finally, $u(x,t)\rightarrow u_{0}(x)$ as $t\rightarrow 0$ by parabolic regularity and compatibility of the initial and boundary data.
Consequently, the function $\psi$ solving the ODE (38) provides a universal stationary parabolic barrier that replaces the mean convexity of the boundary, establishing local–in–time existence of the mean curvature flow of graphs over arbitrary domains.
Remark 11
Let $[0,U)$ denote the maximal time interval of existence of the classical solution $u$ to (39). By the interior estimates of Ecker–Huisken, for any sequence $t_{k}\nearrow U$ there exists a subsequence (still denoted $t_{k}$) such that $u(\cdot,t_{k})$ converges in $C_{\mathrm{loc}}^{\infty}(\Omega)$ to a function $v\in C^{\infty}(\Omega)$. Moreover, $v$ satisfies the stationary minimal surface equation $\mathcal{M}(v)=0$ in $\Omega$, hence defines a minimal graph in the interior of the domain. However, since all these estimates are purely interior, no information can be extracted in general about the behavior of the solution at the boundary as $t\nearrow U$. In particular, the possible obstruction to extending the flow beyond $U$ is entirely of boundary nature.